On a coordinate graph, what's the distance from 6,3 and 4,-6
Answer by CharlesG2(834)

On a coordinate graph, what's the distance from 6,3 and 4,-6
distance D = sqrt((x2-x1)^2+(y2-y1)^2)
D = sqrt((4-6)^2 + (-6-3)^2)
D = sqrt(-2^2 + -9^2)
D = sqrt(4 + 81)
D = sqrt(85) = 9.219544 rounded to 6 places